OpenStudy (mtimko):

A car tire rotates 12.7 rev in 34 s. What is the angular velocity of the tire?

OpenStudy (irishboy123):

\(\large \omega = \frac{2 \pi}{T}\)

OpenStudy (anonymous):

the frequency of rotation , f , is 12.7/34 (Hz). That's the number of revolutions per second. This gives us information about the Time period of rotation , T. T = 2π/w , (where w = angular velocity.) from this we can deduce that w= 2π/T = 2πf. ( because f=1/T ) all that is left is to plug in the numbers: w = 2π (12.7/34) = 2.35 radians/second Ans: 2.35 radians/second

OpenStudy (shamim):

U should not give the answer @AlexandruGuja